# Linzhan Mou5, Linqi Song4, Zhenan Sun [^2] Corresponding authors. 1,3, Ying Wei22footnotemark: 2 6 1

### v2

@@ Appendix G Limitations and Broader Impacts
Our work identifies the presence of massive outliers in down-projection layer of FFN modules, which significantly complicates low-bit weight-activation quantization. To address this challenge, we implement a combination of rotation matrices and permutations to effectively smooth both massive and uniform outliers, proving both fast and effective. Consequently, we establish a new state-of-the-art for INT4 weight-activation post-training quantization methods. Our approach aims to accelerate large language models and reduce memory usage during deployment, offering substantial benefits to the field of LLM research. These advancements could lead to more efficient and accessible LLM applications, facilitating broader usage and enabling more sustainable AI implementations.
Appendix H More Visualizations
We provide additional visualizations of normal and massive outliers in various models (LLaMA1, LLaMA2, Vicuna-v1.5) from Figure H1 to Figure H8. In each figure, the left side illustrates changes in normal outliers before and after applying our rotation and permutation transformations, while the right side shows the changes in massive outliers before and after transformations. It is evident that massive outliers consistently occur in the down-projection layer of the FFN module across all models, supporting our findings discussed in Section 2. Conversely, normal outliers appear in different modules within the transformation block. For instance, Figure H3 shows normal outliers at the up-projection layer of the FFN module in LLaMA1-13B. Significantly, both massive and normal outliers are reduced markedly after our rotation and permutation transformations, leading to easier quantization of activations. This underscores the effectiveness of our RAP in managing outlier features across diverse LLM models.
Figure H1: Activation change with the use of our DuQuant for LLaMA1-7B.
Figure H2: Activation change with the use of our DuQuant for LLaMA1-13B.
Figure H3: Activation change with the use of our DuQuant for LLaMA1-13B.
Figure H4: Activation change with the use of our DuQuant for LLaMA1-65B.
Figure H5: Activation change with the use of our DuQuant for LLaMA2-13B.
Figure H6: Activation change with the use of our DuQuant for LLaMA2-70B.
Figure H7: Activation change with the use of our DuQuant for Vicuna-v1.5-7B.
Figure H8: Activation change with the use of our DuQuant for Vicuna-v1.5-13B.

### v3

@@ Appendix H Limitations and Broader Impacts
Our work identifies the presence of massive outliers in down-projection layer of FFN modules, which significantly complicates low-bit weight-activation quantization. To address this challenge, we implement a combination of rotation and permutation matrices to effectively smooth both massive and uniform outliers, proving both fast and effective. Consequently, we establish a new state-of-the-art for INT4 weight-activation post-training quantization methods. Our approach aims to accelerate large language models and reduce memory usage during deployment, offering substantial benefits to the field of LLM research. These advancements could lead to more efficient and accessible LLM applications, facilitating broader usage and enabling more sustainable AI implementations.
Appendix I More Visualizations
We provide additional visualizations of normal and massive outliers in various models (LLaMA1, LLaMA2, Vicuna-v1.5, Mistral) from Figure I1 to Figure I9. In each figure except Mistral, the left side illustrates changes in normal outliers before and after applying our rotation and permutation transformations, while the right side shows the changes in massive outliers before and after transformations. It is evident that massive outliers consistently occur in the down-projection layer of the FFN module across all models, supporting our findings discussed in Section 2. Conversely, normal outliers appear in different modules within the transformation block. For instance, Figure I3 shows normal outliers at the up-projection layer of the FFN module in LLaMA1-13B. Significantly, both massive and normal outliers are reduced markedly after our rotation and permutation transformations, leading to easier quantization of activations. This underscores the effectiveness of our DuQuant in managing outlier features across diverse LLM models.
Figure I1: Activation change with the use of our DuQuant for LLaMA1-7B.
Figure I2: Activation change with the use of our DuQuant for LLaMA1-13B.
Figure I3: More examples of activation change with the use of our DuQuant for LLaMA1-13B.
Figure I4: Activation change with the use of our DuQuant for LLaMA1-65B.
Figure I5: Activation change with the use of our DuQuant for LLaMA2-13B.
Figure I6: Activation change with the use of our DuQuant for LLaMA2-70B.
Figure I7: Activation change with the use of our DuQuant for Vicuna-v1.5-7B.
Figure I8: Activation change with the use of our DuQuant for Vicuna-v1.5-13B.
Figure I9: Massive activation change with the use of our DuQuant for Mistral7B.